OpenStudy (anonymous):

Solve the equation e^-x=(e^4)^x+5

OpenStudy (anonymous):

1/y = y^4 + 5 (let e^x = y) 1 = y^5 + 5y y^5 + 5y - 1 = 0 y = 0.1999 e^x = 0.1999 x = ln 0.1999 = -1.61

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

Yes thank you so much!